OpenStudy (anonymous):

Simplify each expression and circle your answer. (a^-1 - b^-1) ^-1

OpenStudy (anonymous):

1/ 1/a - 1/b

myininaya (myininaya):

\[(a^{-1}-b^{-1})^{-1}=\frac{1}{a^{-1}-b^{-1}}=\frac{1}{\frac{1}{a}-\frac{1}{b}}=\frac{ab}{ab} \cdot \frac{1}{\frac{1}{a}-\frac{1}{b}}\] \[\frac{ab}{b-a}\]